# Finding radio transients with anomaly detection and active learning based on volunteer classifications

### v1

@@ 5.2 Active Learning Improvement
To quantify how much active learning changes our results, we subtract each updated recall curve from those in section 5.1. The residual curves for all feature-model pairs can be seen in Figure 7. We can immediately see that the wavelet feature instances improve dramatically, increasing recall by as much as 10 per cent within the most anomalous region of feature space. Active learning applied to the feets features improves anomaly recall for the most anomalous sources, but once $\sim$15 per cent of the data have been analysed, minimal improvement is seen. This is reflected in the respective $R_{10}$ and $R_{30}$ values seen in Table 2 which remain almost unchanged to two decimal places. As seen in the previous section, $(\eta,V)$ show no improvement and in some cases perform worse than before. This is in part because our active learning procedure involves labelling the sources at the top of our anomaly score-ordered. Given that the anomaly detection performance using the standard $(\eta,V)$ essentially matches random chance and that transients make up $\sim$2 per cent of our sample, labelling only 2 per cent of the data will not provide the regressor with many positive anomalies from which to learn.
@@
The purity values of Table 2 show minimal improvement following active learning. Only the $P_{10}$ values improve by, at most, 1 per cent, whilst all $P_{30}$ values remain the same to two decimal places. This is in part due to the asymptotic behaviour of these purity calculations as mentioned earlier, namely that as $n$ reaches our total sample size, the purity must approach the overall rate of transients in our data of $\sim$ 2 percent. As a result, it is hard to improve our $P_{30}$ values much when there are so few anomalies to find in our dataset.
@@
In terms of AUC values, we see minimal change compared to the default anomaly detection procedure, with all feets and $(\eta,V)$ values varying at the sub-per cent level and wavelet instances improving by only 0.01. This matches what we see in Figure 7 where after x=0.3 all curves vary only minimally around no change (y=0). As before we stress that this metric does not match the user-driven experience of finding anomalies preferrentially in as small a data volume as possible. As before, the feets-LOF iteration provide the most transients in 10 per cent of the data, though the wavelet instances perform almost as well following the active learning.
@@
Figure 7: Changes in performance due to active learning for each model-feature pair, where $y=0$ indicates no improvement. The x-axis refers to how far down each anomaly score-ordered list is being considered. This plot has been truncated at $x=0.4$ as, after this point, all curves vary minimally about 0, with two figures shown for clarity when comparing between curves. It can be seen that, when using the wavelet features, improvements in recall of up to 12 percentage points can be seen for both anomaly detection algorithms, whilst the feets features improve only moderately
@@ 6.1 Feature performance and comparison
A final comparison between the feets and wavelet features can be seen in Figure 8, where the distributions of actively learnt scores between different classes are shown. In both instances, background (i.e. non-anomalous) sources receive lower anomaly scoures as expected. However, when comparing between previously known transients (45 sources, see A23 for the listed XRBs and other commensal discoveries) and those found by volunteers (142 sources), it can be seen that when using the wavelet features, these two classes produce similar, tightly peaked distributions in score i.e. they are approximately equally anomalous. However, when using the feets features, the distribution of both desired classes are broader, with volunteer-found sources overlapping slightly more with background objects. This implies that, if searching only for sources such as the previously known XRBs, that the feets features may be the feature set to use, but that the wavelet features can find both classes of transient with equal aptitude.
@@
We note that one could spend a great deal longer curating the perfect feature set for a given task and there is no guarantee that the features tested herein form the optimal choices. However, that pre-existing methods work well in this regime is a significant step towards finding radio transients in real-time surveys. Future work may investigate whether, for example, autoencoder architectures (e.g. Villar et al., 2021) provide a way to extract meaningful features directly from the radio time series.
@@ 6.2 Computational Scalability
However, additional steps will be required in order to maximise potential discoveries from current and upcoming data streams. For example, several recent discoveries have been made using fast images (that is, sub-integrations of a single long observation) that warrant further investigation (Hurley-Walker et al., 2022; Caleb et al., 2022; Wang et al., 2023), while in this work the images are typically separated by a week. We plan to apply these anomaly detection methods to sets of 8s-images taken by MeerKAT in the near future. Similarly, we have not made use of any spectral or polarimetric information inherent in the data, both of which can provide crucial additional information and reduce contaminants (Heywood, 2023). However, it is impossible to know on which timescales and at which frequencies to search in a manner that is completely agnostic to all the possible underlying astrophysics. Therefore it is worth thinking about the specific science case in mind and how to best leverage these data products - for example, if one is searching for radio flaring stars, the use of Stokes V imaging is a powerful tool that eliminates many contaminants from images (e.g. Pritchard et al., 2021, 2023). These additional pipeline steps would, however, add to the computational time of finding these anomalies. It is therefore down to the given science case whether data rates are such that only the bare minimum information can be extracted, or whether the astrophysical phenomena necessitate additional processing.
@@ 6.3 Citizen Science
All of the discoveries used as our ground truth were found by volunteers. With upcoming (big) data projects such as the SKA and LSST, it is worth discussing the role of individuals and groups with regards to discovery science. A crude but immediate comparison between human discoveries and automated anomaly detection algorithms can be seen in Figure 9, where the transient vote fraction is how many out of 10 volunteers that saw each source voted for it as a transient/variable. The anomaly ranking is a source’s position in the anomaly score-ordered for the wavelet-LOF instance, though other feature-model pairs were tested with qualitatively similar results. We can see that the majority of volunteer transients are ranked highly by the anomaly detection model, matching the recall curves seen earlier. However, we can also see that there are known transients and variables missed by volunteers (i.e below A23’s threshold of 0.4) - these are typically low significance sources and/or those that show variability that was difficult for volunteers to identify but were found in previous studies i.e. Driessen et al. (2020); Rowlinson et al. (2022); Driessen et al. (2022b). It is encouraging then that the anomaly detection methods rank these anomalies as highly as many of the new transients/variables confirmed by volunteers, indicating that we can avoid a potential source of bias in our searches. Similarly, there are sources given low scores by the anomaly detection algorithms (even with active learning) that were clearly identified by a majority of volunteers - these are the data points about the $y=x$ line in Figure 9. So these are anomalies that volunteers find "more easily" than our algorithms. It is therefore pertinent to leverage each search method in a complementary manner (see also Wright et al., 2017). For example, one could perform a first pass with anomaly detection procedures, which would find the highly anomalous sources easily and then pass the sources that are not ranked highly (e.g. beyond the ‘knee’ of a given recall curve) to citizen scientists. However, it is also important to value the time taken by individual volunteers and consider whether asking them to find relatively few sources in large data volumes is an optimal or even fair division of labour. It is also important to note that this brief comparison with volunteer performance may not necessarily generalise to other projects, future data releases or other volunteers and that project builders could in theory give different training and instructions, producing different results. Nevertheless, the combination of rapid computation and the advantages of volunteer-led searches in concert with expert analysis and, as demonstrated in this work, anomaly detection techniques, hold great promise for future transient hunting efforts.
@@ Acknowledgements
AA acknowledges the support given by the Science and Technology Facilities Council through a STFC studentship. CJL acknowledges support from the Alfred P. Sloan foundation. ML acknowledges support from the South African Radio Astronomy Observatory and the National Research Foundation (NRF) towards this research. Opinions expressed and conclusions arrived at, are those of the authors and are not necessarily to be attributed to the NRF. MV acknowledges financial support from the Inter-University Institute for Data Intensive Astronomy (IDIA), a partnership of the University of Cape Town, the University of Pretoria and the University of the Western Cape, and from the South African Department of Science and Innovation’s National Research Foundation under the ISARP RADIOMAP Joint Research Scheme (DSI-NRF Grant Number 150551) and the CPRR HIPPO Project (DSI-NRF Grant Number SRUG22031677).

### v2

@@ 5.2 Active Learning Improvement
To quantify how much active learning changes our results, we subtract each updated recall curve from those in section 5.1. The residual curves for all feature-model pairs can be seen in Figure 7. We can immediately see that the wavelet feature instances improve increasing recall by as much as 10 per cent within the most anomalous region of feature space. Active learning applied to the feets features improves anomaly recall for the most anomalous sources, but once $\sim$15 per cent of the data have been analysed, minimal improvement is seen. This is reflected in the respective $R_{10}$ and $R_{30}$ values seen in Table 2 which remain almost unchanged to two decimal places. The combined feature set shows a dramatic increase in performance with the LOF algorithm, by up to 20 percentage points at times. This corresponds to recovering approximately half of the transients in 8 per cent of the data compared to only around one third beforehand. By contrast, the combined feature set as applied to the IF shows approximately the same qualitative improvements as when applying the feature sets alone. As seen in the previous section, $(\eta,V)$ show no improvement and in some cases perform worse than before. This is in part because our active learning procedure involves labelling the sources at the top of our anomaly score-ordered list. Given that the anomaly detection performance using the standard $(\eta,V)$ feature pair essentially matches random chance and that transients make up $\sim$2 per cent of our sample, labelling only 2 per cent of the data will not provide the regressor with many positive anomalies from which to learn.
@@
The precision values of Table 2 show minimal improvement following active learning. Only the $P_{10}$ values improve by, at most, 1 per cent, whilst all $P_{30}$ values remain the same to two decimal places. This is in part due to the asymptotic behaviour of these precision calculations as mentioned earlier, namely that as $n$ reaches our total sample size, the precision must approach the overall rate of transients in our data of $\sim$ 2 percent. As a result, it is hard to improve our $P_{30}$ values much when there are so few anomalies to find in our dataset.
@@
In terms of AUC values, we see minimal change compared to the default anomaly detection procedure, with all feets and $(\eta,V)$ values varying at the sub-per cent level and wavelet instances improving by only 0.01. This matches what we see in Figure 7 where after x=0.3 all curves vary only minimally around no change (y=0). As before we stress that this metric does not match the user-driven experience of finding anomalies preferentially in as small a data volume as possible. As before, the feets-LOF iteration provide the most transients in 10 per cent of the data, though the wavelet instances perform almost as well following the active learning.
@@
Figure 7: Changes in performance due to active learning for each feature-model pair, where $y=0$ indicates no improvement. The x-axis refers to how far down each anomaly score-ordered list is being considered. This plot has been truncated at $x=0.4$ as, after this point, all curves vary minimally about 0, with two figures shown for clarity when comparing between curves. It can be seen that, when using the wavelet features, improvements in recall of up to 12 percentage points can be seen for both anomaly detection algorithms, whilst the feets features improve only moderately. When combining feets and wavelet features, there is a 20 percentage point increase when using the LOF, whilst the IF improvement is similar to that of each feature set alone.
@@ 6.1 Feature performance and comparison
The combined feature set performs similar to both individual sets. The recall plot of the combined features is effectively an average of the feets and wavelet feature sets. This implies that while both feature sets contribute information, they are not entirely orthogonal. The additional features will make some sources appear more anomalous while others will now have more in common with the larger population. The base anomaly detection algorithms are not able to make effective use of the additional features. However, applying active learning with the LOF dramatically improves performance of these features, improving recall by up to 20 percentage points. By using this training data, active learning can far more effectively take advantage of the additional information provided, leveraging either the feets features or the wavelets depending on the region of feature space in which the labelled source is found. To avoid these differences caused by the initial anomaly detection algorithm, future active learning procedures may skip this step entirely (Lochner & Rudnick, 2025).
A comparison between the feets and wavelet features can be seen in Figure 8, where the distributions of actively learnt scores between different classes are shown. In both instances, background (i.e. non-anomalous) sources receive lower anomaly scoures as expected. However, when comparing between previously known transients (45 sources, see A23 for the listed XRBs and other commensal discoveries) and those found by volunteers (142 sources), it can be seen that when using the wavelet features, these two classes produce similar, tightly peaked distributions in score i.e. they are approximately equally anomalous. However, when using the feets features, the distribution of both desired classes are broader, with volunteer-found sources overlapping slightly more with background objects. This implies that, if searching only for sources such as the previously known XRBs, that the feets features may be the feature set to use, but that the wavelet features can find both classes of transient with equal aptitude.
@@
We note that one could spend a great deal longer curating the perfect feature set for a given task and there is no guarantee that the features tested herein form the optimal choices. For example, we know that the heterogeneity of our time series and their sampling, along with the imputation of shorter light curves, likely induces biases in the wavelet calculations. However, without labels it would not be possible to tune hyperparameters in any stage of the process. Nevertheless, that pre-existing methods work well in this regime is a significant step towards finding radio transients in real-time surveys. Future work may investigate whether, for example, autoencoder architectures (e.g. Villar et al., 2021) provide a way to extract meaningful features directly from the radio time series.
As a final point on feature selection, we have used Gaussian Processes only for interpolating light curves, though they serve as powerful tools in their own right. When fitting a GP, the covariance kernels used have hyperparameters that can provide alternative metrics for detecting variables and transients. This idea is explored further in Fu et al. (in prep.), who use a multi-term covariance kernel and its hyperparameters to find candidate transients in the same data used here. This study finds that the joint distribution of hyperparameters is a good discriminant for separating variable and transient candidates from stable ones and goes on to detect transients missed by traditional metrics and by A23.
@@ 6.2 Computational Scalability
However, additional steps will be required in order to maximise potential discoveries from current and upcoming data streams. For example, several recent discoveries have been made using fast images (that is, sub-integrations of a single long observation) that warrant further investigation (Hurley-Walker et al., 2022; Caleb et al., 2022; Wang et al., 2023), while in this work the images are typically separated by a week. We plan to apply these anomaly detection methods to sets of 8s-images taken by MeerKAT in the near future. Similarly, we have not made use of any spectral or polarimetric information inherent in the data, both of which can provide crucial additional information and reduce contaminants (Heywood, 2023). However, it is impossible to know on which timescales and at which frequencies to search in a manner that is completely agnostic to all the possible underlying astrophysics. Therefore it is worth thinking about the specific science case in mind and how to best leverage these data products - for example, if one is searching for radio flaring stars, the use of Stokes V imaging is a powerful tool that eliminates many contaminants from images (e.g. Pritchard et al., 2021, 2024). These additional pipeline steps would, however, add to the computational time of finding these anomalies. It is therefore down to the given science case whether data rates are such that only the bare minimum information can be extracted, or whether the astrophysical phenomena necessitate additional processing.
@@ 6.3 Citizen Science
All of the discoveries used as our ground truth were found by volunteers. With upcoming (big) data projects such as the SKA and LSST, it is worth discussing the role of individuals and groups with regards to discovery science. A crude but immediate comparison between human discoveries and automated anomaly detection algorithms can be seen in Figure 9, where the transient vote fraction is how many out of 10 volunteers that saw each source voted for it as a transient/variable. The anomaly ranking is a source’s position in the anomaly score-ordered list for the wavelet-LOF instance, though other feature-model pairs were tested with qualitatively similar results. We can see that the majority of volunteer transients are ranked highly by the anomaly detection model, matching the recall curves seen earlier. However, we can also see that there are known transients and variables missed by volunteers (i.e below A23’s threshold of 0.4) - these are typically low significance sources and/or those that show variability that was difficult for volunteers to identify but were found in previous studies i.e. Driessen et al. (2020); Rowlinson et al. (2022); Driessen et al. (2022b). It is encouraging then that the anomaly detection methods rank these anomalies as highly as many of the new transients/variables confirmed by volunteers, indicating that we can avoid a potential source of bias in our searches. Similarly, there are sources given low scores by the anomaly detection algorithms (even with active learning) that were clearly identified by a majority of volunteers - these are the data points about the $y=x$ line in Figure 9. So these are anomalies that volunteers find "more easily" than our algorithms. It is therefore pertinent to leverage each search method in a complementary manner (see also Wright et al., 2017). For example, one could perform a first pass with anomaly detection procedures, which would find the highly anomalous sources easily and then pass the sources that are not ranked highly (e.g. beyond the ‘knee’ of a given recall curve) to citizen scientists. However, it is also important to value the time taken by individual volunteers and consider whether asking them to find relatively few sources in large data volumes is an optimal or even fair division of labour. We caveat that this brief comparison with volunteer performance may not necessarily generalise to other projects, future data releases or other volunteers and that project builders could in theory give different training and instructions, producing different results. Nevertheless, the combination of rapid computation and the advantages of volunteer-led searches in concert with expert analysis and, as demonstrated in this work, anomaly detection techniques, hold great promise for future transient hunting efforts.
Involving citizen scientists at the active learning stage is one way to leverage this combination of algorithmic efficiency and volunteer diligence. As simulated and then demonstrated by the Galaxy Zoo team (Walmsley et al., 2020, 2022a), volunteers can be preferentially shown sources which are most informative for a model to train on, dramatically reducing the number of votes required overall. In an unsupervised setting this human-in-the-loop cycle might constitute showing volunteers the most anomalous sources after default anomaly detection and then updating an acquisition function based on the citizen scientists’ classifications. Volunteers could then be shown sources according to the new anomaly score-ordered list and the process repeats. Due to the subjective nature of what a given (citizen) scientist might find interesting, the resulting acquisition function would be some average of many different opinions, weighted by the number of sources classified per volunteer. Project scientists would then have to decide whether or not to intentionally attempt to influence behaviour by, for example, displaying ’wanted’ source classes in volunteer training, or whether a diversity in voting patterns can be leveraged for downstream tasks. In any case, optimising what sources are shown to volunteers will be of paramount importance when dealing with future surveys such as from the SKA or Vera C. Rubin Observatory.
@@ Acknowledgements
AA acknowledges the support given by the Science and Technology Facilities Council through a STFC studentship and the support of the Breakthrough Listen project. Breakthrough Listen is managed by the Breakthrough Initiatives, sponsored by the Breakthrough Prize Foundation. CJL acknowledges support from the Alfred P. Sloan foundation. ML acknowledges support from the South African Radio Astronomy Observatory and the National Research Foundation (NRF) towards this research. Opinions expressed and conclusions arrived at, are those of the authors and are not necessarily to be attributed to the NRF. JvdE acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation, and was supported by funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101148693 (MeerSHOCKS) for part of this work. MV acknowledges financial support from the Inter-University Institute for Data Intensive Astronomy (IDIA), a partnership of the University of Cape Town, the University of Pretoria and the University of the Western Cape, and from the South African Department of Science and Innovation’s National Research Foundation under the ISARP RADIOMAP Joint Research Scheme (DSI-NRF Grant Number 150551) and the CPRR HIPPO Project (DSI-NRF Grant Number SRUG22031677).